Consider the demand for a good. At price Rs 4, the demand for the good is 25 units. Suppose price of the good increases to Rs 5, and as a result, the demand for the good falls to 20 units. Calculate the price elasticity?


1
Expert's answer
2021-10-06T13:07:00-0400

Price Elasticity of demand is given by percentage change in quantity divided by percentage change in price.

Percentage change in quantity"= \frac{Q_2-Q_1}{(Q_2+Q_1)/2}\times100"


"= \frac {20-25}{(20+25)/2}\times100"


"=-22.22" %

Percentage change in price:"=\frac{ P_2-P_1}{(P_2+P_1)/2}\times100"


"=\frac{5-4}{(5+4)/2}\times100"

"=22.22" %

Price Elasticity "= \frac{-22.22}{22.22}= -1"